# FaceCloak: Learning to Protect Face Templates

### v1

@@ Abstract
Generative models can reconstruct face images from encoded representations (templates) bearing remarkable likeness to the original face raising security and privacy concerns. We present FaceCloak, a neural network framework that protects face templates by generating smart, renewable binary cloaks. Our method proactively thwarts inversion attacks by cloaking face templates with unique disruptors synthesized from a single face template on the fly while provably retaining biometric utility and unlinkability. Our cloaked templates can suppress sensitive attributes while generalizing to novel feature extraction schemes and outperforms leading baselines in terms of biometric matching and resiliency to reconstruction attacks. FaceCloak-based matching is extremely fast (inference time cost=0.28ms) and light-weight (0.57MB).
@@ I INTRODUCTION
Motivation. Face recognition systems (FRS) are susceptible to various forms of attacks from motivated adversaries [17, 18, 29]. Template inversion attacks attempt to invert a given face template, typically produced using a deep neural feature extraction model, to recover its corresponding original face image [12, 28, 4, 10]. In order to carry out such an attack, the adversary requires access to the target template, and either white-box or black-box access to the feature extractor model. An inverted template can pose multiple concerns: the adversary can then use the original biometric sample to generate system-specific templates for unauthorized access to multiple services (banking, phone, medical records etc.). Moreover, the adversary can deduce protected attributes such as gender, age, and ethnicity of the target, raising privacy concerns. Some well-known face template inversion methods that use generative models are NbNet, StyleGAN-inversion and Arc2Face [28, 32, 34]. Template-level attacks are not new, but lately, novel attacks are emerging with higher diversity and increased attack success due to three factors: (i) rapid surge of IoT devices connecting users to remote applications (mobile banking, online interview); (ii) widespread deployment of biometric authentication for easy access (Apple FaceID, Google Pay using Face Unlock); (iii) a gamut of open-source generative tools that can instantly simulate realistic looking synthetic media, thus attracting the attention of malicious agents (diffusion models, LLMs).
As a result, biometric template protection schemes are a recommended security measure in real-world FRS deployments [24]. Most template protection schemes are based on either semantic feature transformations of the given template [3, 35, 15], or cryptographic encryption schemes [26, 20, 11, 5], or hybrid approaches. However, all such schemes suffer from performance-security trade-offs or can leak sensitive attributes [42].
@@
FaceCloak uses a neural network that takes a single unprotected face template as input, then derives a set of disruptors consisting of positive (noise and mask), and negative instances (synthetic templates and orthogonalization); finally combines the original template with the disruptors to construct the protected or cloaked template supervised by biometric identity, binarization and diversity losses. Our network with two fully connected layers, batch normalization and activations is non-trivial to invert while the loss functions retain matching performance; see Fig 1. During enrollment, the network learns to optimally combine disruptors with the original template. After training, we keep the trained network and the protected gallery template while discarding the disruptors. During verification, the query template goes through a single pass of the trained network, and is compared with the enrolled protected template using Hamming distance. Our method does not require user-specific key management, is efficient and highly effective in ensuring biometric performance, security and privacy.
@@
Our method is generalizable to novel feature extraction schemes. For validation, we propose ArcFace-OPL (ArcFace finetuned with Orthogonal Projection Loss) for improving disruptor creation and test FaceCloak for robustness. We also demonstrate that our protected templates provide user privacy by suppressing sensitive demographic attributes, such as, gender.
@@ II PRIOR WORK
Face template protection comes under the purview of biometric template protection (BTP) which focuses on protecting the biometric templates to safeguard the privacy of the owner and at the same time provide security to the authentication framework. They focus on ensuring irreversibility of the protected templates, i.e., the protected templates will be resilient against inversion attacks, retention of biometric performance, i.e., minimal loss of biometric utility, and unlinkability, i.e., a new protected template can be regenerated which will be unlinked to the previously compromised protected template. Classical approaches apply feature transformation to convert original face templates to protected templates with the help of user-specific transformation functions [3, 35, 15]. Homomorphic encryption (HE) is one of the most popular face BTP schemes as it allows comparison in the encrypted domain while ideally causing no biometric performance degradation. However, computational overhead and the need for maintaining the decryption algorithm as secret pose disadvantages for HE. Currently, there has been significant progress in speeding up HE operations through quantization and dimensionality reduction [26, 5, 11, 20]. Fuzzy Commitment scheme [31, 13] and Fuzzy Vault scheme [21] apply cryptographic hash functions to randomly generated codewords, which are related to the original face templates via a mathematical function. Secure sketch for biometric templates utilized auxiliary user information and provides a bound on the performance loss in terms of relative entropy [25]. The mapping between the face template and the codeword can be established via the Locality Sensitive Hashing (LSH) algorithm [7], while IronMask [23] defined the mapping in terms of a linear transformation represented by an orthogonal matrix. Secure Face [27] uses a randomized CNN and user-specific key for template protection. A deep rank hashing (DRH) network and a cancellable identification scheme was proposed in [9]. BioHashing [40] performed template protection via discretization of original template using a seed value and random number, and a secret key sharing scheme. MLP-hash [39] used a user-specific randomly weighted MLP to create binary protected templates. Index-of-Maximum (IoM) [19] hashing used ranking of LSH for template protection. Refer to [24] for a comprehensive survey. Recently, SLERPFace [43] used spherical and linear interpolation of templates to rotate on the hypersphere to become noise. Existing methods rely on user-specific parameters/keys [15, 5] or computationally prohibitive [39, 23] or is restricted to a specific type of templates like SLERPFace [43] requires unit normalized features based on angular margin but does not analyze generalizability across other features such as FaceNet [37] or MagFace [30].
@@ III PROPOSED METHOD
Biometric identity loss: We use triplet learning with margin, (hyperparameter) and MSE loss using the original template as anchor (a), positive disruptors (p), and negative disruptors (n).
@@ IV EXPERIMENTS AND ANALYSIS
| Metric | Description | ArcFace | FaceNet | ||
| Biometric Matching |
96.2% | 69.3% | ||
95.7% | 67.8% | |||
| Irreversibility | SAR | 0.0% | 0.0% | ||
| Unlinkability | 0.003 | 0.03 | |||
@@
Generalizability. We rely on the face template extracted using existing schemes such as ArcFace and FaceNet to derive positive and negative disruptors. Therefore, the selection and creation of optimal disruptors is important for driving FaceCloak. So, we investigated a secondary supervision strategy that uses Orthogonal Projection Loss (OPL) [36] in conjunction with angular margin loss () and performed preliminary experiments on the LFW dataset to test the generalizability of our method. OPL is formulated as: Here, denotes cosine similarity operator, denotes normalized features/templates belonging to classes , and denotes absolute operator. We trained the ArcFace model (IResNet 50) [2] package on the VGGFace2 [6] dataset (200K images 8K identities) with the new loss , where . We test the robustness of FaceCloak using unprotected templates derived using on the LFW dataset. We achieve TMR=90.1% @ FMR=0 after cloaking and unlinkability values as =0.01 and =0.05 which shows promising results in terms of (i) capability of OPL for effective disruptor creation and (ii) robustness of FaceCloak towards novel loss functions for face template extraction. We will investigate its potential further in future work.
Demographic attribute protection. PolyProtect can leak sensitive attributes [42] and need homomorphic encryption for user privacy. We test how FaceCloak performs when a neural network tries to extract demographic (gender) information from our cloaks (protected templates). We first train a 2-layer fully connected network (64321) followed by ReLU and sigmoid activations, respectively. We train the network on a subset of cloaks generated from the LFW dataset (3,200) and test on a disjoint set of cloaks (1,400) such that they are balanced in terms of gender distribution. We use DeepFace to extract ground-truth gender labels. We train the network using Adam optimization for 100 epochs and use Binary Cross Entropy loss. We achieve 48.4% gender prediction accuracy (close to random chance=50%). This shows that our method is capable of suppressing sensitive demographic attributes successfully.
@@ V SUMMARY AND FUTURE WORK
We design a neural network driven novel face template protection scheme known as FaceCloak that firstly, learns an optimal set of disruptors from a single unprotected face template, and secondly, combines them with the face template to create a binary cloak. FaceCloak acts as a hash function and is supervised via biometric identity, binarization and diversity losses. The novelty of our method lies in leveraging a small set of easily regenerated disruptors that guide the network in retaining face matching while mitigating inversion attacks. Experiments show that our method achieves irreversibility, biometric performance, and unlinkability/renewability while eliminating the need for secret key or user-specific parameter management, and can be easily deployed as an add-on module with minimal overhead. Future work will further test the robustness of our method across diverse datasets, feature extraction schemes, modalities, and against attack via record multiplicity (ARM), and additional baselines.
@@
We have developed a face template protection scheme that protects the privacy of an individual by mitigating template inversion attacks. It further secures the biometric recognition system. Our method can be used as an additional module in existing systems with little overhead. Our algorithm uses synthetic faces generated using StyleGAN3. As the generative model was trained on real faces, it may lead to inadvertent information leakage. To avoid this, we ensure that the disruptors (which use synthetic faces) are never stored and can be created on the fly. We test the security of the proposed method by considering a white-box attack and our preliminary analysis shows strong resilience.

### v2

@@ Abstract
Generative models can reconstruct face images from encoded representations (templates) bearing remarkable likeness to the original face, raising security and privacy concerns. We present FaceCloak, a neural network framework that protects face templates by generating smart, renewable binary cloaks. Our method proactively thwarts inversion attacks by cloaking face templates with unique disruptors synthesized from a single face template on the fly while provably retaining biometric utility and unlinkability. Our cloaked templates can suppress sensitive attributes while generalizing to novel feature extraction schemes and outperform leading baselines in terms of biometric matching and resiliency to reconstruction attacks. FaceCloak-based matching is extremely fast (inference time =0.28 ms) and light (0.57 MB). We have released our code for reproducible research.
@@ I INTRODUCTION
Motivation. Face recognition systems (FRS) are susceptible to various forms of attacks by motivated adversaries [17, 18, 29]. Template inversion attacks attempt to invert a given face template, typically produced using a deep neural feature extraction model, to recover its original face image [12, 28, 4, 10]. In order to carry out such an attack, the adversary requires access to the target template and either white-box or black-box access to the feature extractor model. An inverted template can pose multiple concerns: the adversary can then use the original biometric sample to generate system-specific templates for unauthorized access to multiple services (banking, phone, medical records, etc.). In addition, the adversary can deduce protected attributes such as the gender, age, and ethnicity of the target, raising privacy concerns. Some well-known face template inversion methods that use generative models are NbNet, StyleGAN-inversion and Arc2Face [28, 32, 34]. Template-level attacks are not new, but lately, novel attacks are emerging with higher diversity and increased attack success due to three factors: (i) rapid surge of IoT devices connecting users to remote applications (mobile banking, online interview); (ii) widespread deployment of biometric authentication for easy access (Apple FaceID, Google Pay using Face Unlock); (iii) a gamut of open-source generative tools that can instantly simulate realistic looking synthetic media, thus attracting the attention of malicious agents (diffusion models, LLMs).
As a result, biometric template protection schemes are recommended security measures in real-world FRS deployments [24]. Most template protection schemes are based on semantic feature transformations of the given template [3, 35, 15], or cryptographic encryption schemes [26, 20, 11, 5], or hybrid approaches. However, all such schemes suffer from performance-security trade-offs or can leak sensitive attributes [42].
@@
FaceCloak uses a neural network that takes a single unprotected face template as input, then derives a set of disruptors consisting of positive (noise and mask), and negative instances (synthetic templates and orthogonalization); finally combines the original template with the disruptors to construct the protected or cloaked template supervised by biometric identity, binarization and diversity losses. Our network with two fully connected layers, batch normalization and activations is non-trivial to invert while the loss functions retain matching performance; see Fig 1.
Use case: During enrollment, FaceCloak accepts any face template as input, and generates disruptors to create a protected template. The gallery will save the trained network and its output as the enrolled template, but discard the disruptors. During verification, the query template goes through the trained FaceCloak. The output is then compared with the enrolled protected template from the gallery using Hamming distance for decision. Our method does not require user-specific key management, is efficient and highly effective in ensuring biometric matching, security and privacy.
@@
Our method is generalizable to novel feature extraction schemes. For validation, we propose ArcFace-OPL (ArcFace finetuned with Orthogonal Projection Loss) for improving disruptor creation and test FaceCloak for robustness. We further demonstrate that our template protection scheme can inherently suppress sensitive demographic attributes, such as, gender.
@@ II PRIOR WORK
Face template protection comes under the purview of biometric template protection (BTP) which focuses on protecting the biometric templates to safeguard the privacy of the owner and at the same time provide security to the authentication framework. They focus on ensuring irreversibility of the protected templates, i.e., the protected templates will be resilient against inversion attacks, retention of biometric performance, i.e., minimal loss of biometric utility, and unlinkability, i.e., a new protected template can be regenerated which will be unlinked to the previously compromised protected template. Classical approaches apply feature transformation to convert original face templates to protected templates with the help of user-specific transformation functions [3, 35, 15]. Homomorphic encryption (HE) is one of the most popular face BTP schemes as it allows comparison in the encrypted domain while ideally causing no biometric performance degradation. However, computational overhead and the need for maintaining the decryption algorithm as secret pose disadvantages for HE. Currently, there has been significant progress in speeding up HE operations through quantization and dimensionality reduction [26, 5, 11, 20]. Fuzzy Commitment scheme [31, 13] and Fuzzy Vault scheme [21] apply cryptographic hash functions to randomly generated codewords, which are related to the original face templates via a mathematical function. Secure sketch for biometric templates utilized auxiliary user information and provides a bound on the performance loss in terms of relative entropy [25]. The mapping between the face template and the codeword can be established via the Locality Sensitive Hashing (LSH) algorithm [7], while IronMask [23] defined the mapping in terms of a linear transformation represented by an orthogonal matrix. Secure Face [27] uses a randomized CNN and user-specific key for template protection. A deep rank hashing (DRH) network and a cancellable identification scheme was proposed in [9]. BioHashing [40] performed template protection via discretization of original template using a seed value and random number, and a secret key sharing scheme. MLP-hash [39] used a user-specific randomly weighted MLP to create binary protected templates. Index-of-Maximum (IoM) [19] hashing used ranking of LSH for template protection. Refer to [24] for a comprehensive survey. Recently, SLERPFace [43] used spherical and linear interpolation of templates to rotate on the hypersphere to become noise. Existing methods either rely on user-specific parameters/keys [15, 5], or can be computationally prohibitive [39, 23] or maybe restricted to a specific type of feature extractor, such as SLERPFace [43], which requires unit normalized features based on angular margin, and does not analyze generalizability across other features such as FaceNet [37] or MagFace [30].
@@ III PROPOSED METHOD
Biometric identity loss: We use triplet margin with distance loss with the original template as anchor (a), positive disruptors (p), negative disruptors (n), and a tunable margin parameter, .
@@ IV EXPERIMENTS AND ANALYSIS
| Metric | Description | ArcFace | FaceNet | ||||
96.2% | 69.3% | ||||
95.7% | 67.8% | |||||
| Irreversibility | SAR | 0.0% | 0.0% | ||||
| Unlinkability | 0.003 | 0.03 | |||||
@@
Generalizability. We rely on the face template extracted using existing schemes such as ArcFace and FaceNet to derive positive and negative disruptors. Therefore, the selection and creation of optimal disruptors is important for driving FaceCloak. So, we investigated a secondary supervision strategy that uses Orthogonal Projection Loss (OPL) [36] in conjunction with angular margin loss () and performed preliminary experiments on the LFW dataset to test the generalizability of our method. OPL is formulated as: Here, denotes cosine similarity operator, denotes normalized features/templates belonging to classes , and denotes absolute operator. We trained the ArcFace model (IResNet 50) [2] package on the VGGFace2 [6] dataset (200K images 8K identities) with the new loss , where . We tested the robustness of FaceCloak using unprotected templates derived using on the LFW dataset. We achieve TMR=90.1% @ FMR=0 after cloaking and unlinkability values as =0.01 and =0.05 which shows promising results in terms of (i) capability of OPL for effective disruptor creation and (ii) robustness of FaceCloak towards novel loss functions for face template extraction. We will investigate its potential further in future work.
Demographic attribute protection. PolyProtect can leak sensitive attributes [42] and require homomorphic encryption for user privacy. We test whether FaceCloak-based protected templates divulge demographic (gender) information. Note that FaceCloak is designed to be a template protection scheme, and not as a de-identification method for suppressing soft biometrics. We first train a 2-layer fully connected network (64321) followed by ReLU and sigmoid activations, respectively. We train the network on a subset of cloaks generated from the LFW dataset (3,200) and test on a disjoint set of cloaks (1,400) such that they are evenly distributed across gender. We use DeepFace to extract ground-truth gender labels, and train the network using Adam optimization for 100 epochs with Binary Cross Entropy loss. We achieve 48.4% gender prediction accuracy (close to random chance=50%), showing, our capability of of successfully suppressing sensitive demographic attributes.
@@ V SUMMARY AND FUTURE WORK
We design a neural network-driven novel face template protection scheme known as FaceCloak that first learns an optimal set of disruptors from a single unprotected face template and then combines them with the face template to create a binary cloak. FaceCloak acts as a hash function and is supervised by biometric identity, binarization and diversity losses. The novelty of our method lies in the ability to take advantage of a small set of easily regenerated disruptors that guide the network to retain face matching while mitigating inversion attacks. Experiments show that our method achieves irreversibility, biometric performance, and unlinkability/renewability while eliminating the need for secret key or user-specific parameter management, and can be easily deployed as an add-on module with minimal overhead.
Future work will focus on evaluating our method across complex datasets, additional template extractors, modalities, and robustness to attack via record multiplicity (ARM).
@@
We have developed a face template protection scheme that protects the biometric signature of an individual by mitigating template inversion attacks, and further secures the biometric system. Our method can be used as an additional module in existing systems with little overhead. Our algorithm uses synthetic faces generated using StyleGAN3. As the generative model was trained on real faces, it may lead to inadvertent information leakage. To avoid this, we ensure that the disruptors from synthetic faces can be created on the fly. Our preliminary analysis indicates resilience against white-box attack.